### v1

# GLIMPSE-D: An Exotic Balmer-Jump Object at ? Revisiting Photometric Selection and the Cosmic Abundance of PopĀ III Galaxies

Abstract
We present deep JWST/NIRSpec G395M spectroscopy of GLIMPSE-16043, a promising PopĀ III candidate originally identified through NIRCam photometry as having weak [Oāiii] emission. Our follow-up reveals clear [Oāiii] emission, ruling out a genuine zero-metallicity nature. However, the combination of the measured line fluxes and photometry indicates that its spectral energy distribution requires an extraordinarily strong Balmer jump ( mag) and H equivalent width (āĆ ), features that cannot be reproduced by current stellar+nebular or pure nebular photoionization models. The only models approaching the observations to almost within involve a hot ( K) single blackbody embedded in a low- nebular environment, suggestive of scenarios such as a tidal-disruption event or a microquasar with strong disk winds. This cautions that photometric PopĀ III selections are vulnerable to contamination when the rest-frame optical continuum is undetected. Motivated by this, we refine the photometric PopĀ III selection criteria to exclude the locus of extreme Balmer-jump objects. The revised criteria also recover the recently reported spectroscopic candidate AMORE6, demonstrating that the updated selection preserves sensitivity to genuine PopĀ IIIālike sources while removing key contaminants. Applying the refined criteria across legacy survey fields and five newly released CANUCS lensing cluster fields, we revisit the PopĀ III UV luminosity function and estimate the PopĀ III cosmic star-formation rate density to be āĀ Ā yr-1Ā cMpc-3 at ā7, falling in the range of current theoretical predictions.
1 Introduction
The formation of the first stars, known as PopulationĀ III (PopĀ III), marks a pivotal epoch in cosmic history. Emerging from pristine, metal-free gas, these stars initiated the enrichment of the interstellar medium (ISM) and enabled the subsequent formation of chemically evolved galaxies, including the Milky Way (e.g., bromm2013; frebel2015; klessen2023). Their massive nature, strong ionizing spectra, and short lifetimes make PopĀ III stars critical to understanding early structure formation, cosmic reionization, and the seeding of the first black holes (e.g., tumlinson2006; bromm2011; xu2016; inayoshi2020).
Despite their theoretical importance, PopĀ III stars and their host galaxies have remained elusive. Prior to the launch of the James Webb Space Telescope (JWST), searches for PopĀ III signatures were limited to rest-UV diagnostics in relatively bright systems (e.g., nagao2008; sobral2015; shibuya2018), which may not probe the faintest, most metal-poor regimes where PopĀ III formation is expected to occur (e.g., bromm2013). The advent of JWST has revolutionized the search by enabling access to ultra-faint galaxies at with sufficient sensitivity to detect rest-frame optical emission lines, such as H and [OĀ iii], that are key to inferring metallicity and stellar population properties (e.g., chemerynska2024b).
In this context, fujimoto2025 recently presented a novel photometric selection method to identify PopĀ III galaxy candidates at ā7 using deep JWST/NIRCam imaging (see also e.g., nagao2008; inoue2011; zackrisson2011; trussler2023; nishigaki2023). Their search across Ā arcmin2 of legacy survey data led to the discovery of a highly promising candidate, GLIMPSE-16043 (hereafter GLIMPSE-16043), at . This object exhibited key signatures of a PopĀ III-dominated stellar population based on one of the deepest NIRCam photometry so far obtained: extremely strong H emission, a pronounced Balmer jump, an absence of detectable [OĀ iii] emission, a blue UV slope, and an inferred stellar age of Ā Myr. These properties placed ID16043 well within the fiducial PopĀ III selection window and motivated the need for spectroscopic follow-up to confirm its nature.
In this paper, we present the results from a dedicated JWST/NIRSpec deep follow-up program targeting GLIMPSE-16043. The program, dubbed GLIMPSE-D (PID 9223; PIs: S.Ā Fujimoto & R.Ā Naidu), obtained deep NIRSpec/MSA G395M observations to directly assess the presence or absence of key diagnostic emission lines at rest-frame optical wavelengths. As we discuss in SectionĀ 3, our spectroscopic analysis concludes that GLIMPSE-16043 exhibits clear [OĀ iii] emission, definitively ruling out a zero-metallicity PopĀ III origin. Nonetheless, GLIMPSE-16043 turns out to be exceptionally unusual from a different aspect: its Balmer jump and the equivalent widths of the Balmer lines are remarkably high, which cannot be reproduced by current general stellar population and photoionization models. We discuss how such objects may mimic PopĀ III photometric colors, and we reassess the robustness and limitations of current photometric selection techniques.
This paper is structured as follows. In SectionĀ 2, we describe the observations and data reduction. SectionĀ 3 outlines the analysis procedures, including emission line measurements and spectral modeling. SectionĀ 4 presents the main spectroscopic results and investigates the physical origin of GLIMPSE-16043. In SectionĀ 5, we re-evaluate the photometric PopĀ III selection method in light of these findings and provide revised constraints on the PopĀ III UV luminosity function and cosmic star formation rate density. We summarize our conclusions in SectionĀ 6.
Throughout this paper, we adopt a flat CDM cosmology with , , Ā kmĀ s-1Ā Mpc-1. Magnitudes are given in the AB system (oke1983).
2 Observations and Data Reduction
The deep spectroscopic follow-up observations of GLIMPSE-16043 were carried out with JWST/NIRSpec as part of the Directorās Discretionary Time (DDT) program GLIMPSE-D (PID 9223; PIs: S.Ā Fujimoto & R.P.Ā Naidu). The observations were designed to confirm the PopĀ III nature of this photometrically selected candidate by directly testing for the presence of strong H and H emission and the absence of significant [OĀ iii] emission. The data were collected using the NIRSpec/Multi-Object Spectroscopy (MOS) mode with the G395M grating and F290LP filter, covering the wavelength range mām at spectral resolution.
Three separate micro-shutter array (MSA) configurations were used to maximize the number of filler targets while acquiring the full exposure for the main target. These configurations are referred to internally as glimpse_02, glimpse_01, and glimpse_01b, each observed for 42,016Ā sec, 33,613Ā sec, and 33,613Ā sec, respectively. The observations were conducted in two visits: one mask on UT 2025 June 30 and two masks on UT 2025 July 2, resulting in a total on-source integration time of 109,242Ā s (Ā hr) with a consistent MSA position angle. To fully utilize the multiplexing capability of the MSA, additional filler slits were assigned to galaxies with photometric redshifts , enabling ancillary science on faint, lensed galaxies in the AbellĀ S1063 field (see also, e.g., fei2025; kokorev2025b; berg2025).
The data reduction was performed using the msaexp pipeline (brammer2023), following the procedures described in heintz2024 and degraaff2025a. The reduction methodology is identical to that adopted in naidu2025 and in the upcoming v4 release of the DAWN JWST NIRSpec archive (Pollock et al., in prep.), where the flux calibration has been updated. In FigureĀ 1, we show the reduced 2D spectra obtained in the three exposures and the shutter configurations around GLIMPSE-16043 overlaid on the NIRCam F444W image cutout.
3 Data analysis
3.1 Co-adding Spectra
The NIRSpec/G395M observations of GLIMPSE-16043 were obtained in three separate MSA configurations. We first scaled each spectrum to a common absolute flux using the NIRCam F444W photometry as a reference, as this band is well within the G395M wavelength coverage and detected with the best S/N owing to the flux boosting of H line. The synthetic F444W photometry is computed from each reduced spectrum with the F444W filter response curve, and compared to the measured F444W photometric flux to derive a scaling factor. The scaled spectra are then interpolated onto a common wavelength grid and co-added using weights proportional to the H S/N in each configuration. The co-added flux is taken as the weighted mean of the individual spectra, with uncertainties from the inverse-variance sum. In the bottom panel of FigureĀ 1, we present the co-added spectrum, showing the clear detection of H, [Oāiii], and H lines, with no detection of [Heāii]4686. We use this co-added spectrum for all subsequent line measurements throughout this paper.
3.2 Line Flux Measurements
We measured the fluxes of the H, [Oāiii]Ā 4960,5008, and H emission lines from the co-added 1D spectrum by fitting Gaussian profiles. For H and H, we used single-Gaussian models. For [Oāiii], we fit a double-Gaussian model with the flux ratio of 5008 to 4960 fixed to the theoretical value of 2.98, enforcing a common redshift and width for the two components. The fits were performed in narrow wavelength windows around the expected line centers. Thus, we adopt a constant continuum as a free parameter. Because we find that the redshift forced to be exactly the same number makes the fitting worse, we leave the redshift as a free parameter among these three line fitting.
In the inset panels at the bottom panel of FigureĀ 1, we show the best-fit Gaussian for the three emission lines. The individual line redshifts are consistent within the uncertainties (Ā kmĀ s-1), and we obtain the best-fit spectroscopic redshift of GLIMPSE-16043 to be . The resulting line fluxes, S/N, and key flux ratios are summarized in TableĀ 1. Note that we did not identify the [Heāii]4686 line with in the same Gaussian fitting procedure as described above. Using the measurement uncertainty of the H line flux, we also include a 3 upper limit of [Heāii]4686 in TableĀ 1.
| Line | S/N | Flux |
| [Heāii]4686 | ||
| H | 11.2 | |
| Oāiii | 22.7 | |
| H | 23.3 | |
| H/H | ā | |
| Oāiii/H | ā |
4 True Picture of GLIMPSE-16043
4.1 Possible Causes
The clear detection of the [Oāiii] emission line in our G395M spectrum shows that GLIMPSE-16043 is not a truly metal-free (PopĀ III) galaxy. In fujimoto2025, the classification of GLIMPSE-16043 as a PopĀ III candidate was based primarily on the lack of a significant flux excess in the F356W band, where the [Oāiii] line would be expected to fall given its photometric redshift. Given the fact that GLIMPSE-16043 turns out not to be a PopĀ III galaxy, the possible explanations can be divided into four scenarios: (i) an incorrect redshift, (ii) photometric uncertainties, (iii) non-zero dust attenuation (which would suppress H relative to H), or (iv) an exotic Balmer jump. Below, we explain each scenario and discuss whether it is plausible.
(i) Incorrect redshift ā The PopĀ III colorāselection method in fujimoto2025 exploits the flux excesses in F356W and F444W, relative to F410M, arising from [Oāiii]+H and H, respectively. This technique is explicitly valid only for galaxies at ā. The spectroscopic redshift from G395M, , falls well within this range, and the [Oāiii] line is indeed covered by F356W. Thus, an incorrect redshift is not the cause.
(ii) Photometric uncertainties ā From the photometric flux-excess measurements, fujimoto2025 reported a upper limit of [Oāiii]/H. Our G395M measurements yield [Oāiii]/H, a discrepancy with respect to the earlier limit. Such a large difference is unlikely to be explained solely by photometric uncertainties.
(iii) Non-zero dust attenuation ā In fujimoto2025, the contribution of H to the F356W flux excess was subtracted assuming an intrinsic ratio of H/H (i.e., caseĀ B recombination without dust), based on the rest-frame UV SED shape. If significant dust attenuation were present, H would be weaker than assumed, increasing the fraction of the F356W excess attributable to [Oāiii]. However, our observed ratio of H/H is consistent with little to no dust attenuation assuming CaseĀ B recombination, which makes this explanation unlikely.
(iv) Exotic Balmer jump ā Measuring the F356W excess requires a reference continuum filter. In the GLIMPSE NIRCam configuration, F410M is expected to sample the underlying continuum of [Oāiii]+H and H for ā6.6. Notably, GLIMPSE-16043 is undetected in F410M, and in fujimoto2025 the continuum level was estimated from the best-fit PopĀ III SED model to infer the F356W and F444W excesses. If the true continuum is substantially fainter than predicted, the F356W excess would be correspondingly larger, leaving more room for [Oāiii] to contribute. Interestingly, the best-fit PopĀ III SED in fujimoto2025 already features a strong Balmer jump and strong rest-frame H equivalent width (EW) of to satisfy the stringent F410M limit, and requiring an even fainter continuum implies an unusually large Balmer jump and EW(H) in GLIMPSE-16043.
In summary, among scenarios (i)ā(iv), only (iv) an exceptionally strong Balmer jump remains plausible. In the following, we use the new G395M data to quantify the underlying continuum level, the Balmer jump strength, and the implied nebular conditions in GLIMPSE-16043.
4.2 Measuring Continuum and Balmer Jump Strength
In the NIRCam photometry presented in fujimoto2025, the rest-frame optical continuum is constrained only by upper limits for GLIMPSE-16043. With the G395M spectrum now providing measurements of the key rest-frame optical line fluxes, we can estimate the continuum by subtracting the line contributions from the relevant photometric bands. However, the continuum of GLIMPSE-16043 is extremely faint, and in the fluxāscaling procedure described in SectionĀ 3.1, its contribution to the total photometric flux is negligible. As a result, small uncertainties in the absolute flux scaling could have a disproportionate impact on a direct continuum estimate. To mitigate this, we estimate the underlying continuum using the following approach.
First, we adopt the observed [Oāiii]/H ratio of , which is minimally impacted by potential uncertainties in the absolute flux calibration. Second, we fix the intrinsic caseĀ B Balmer decrement to H/H and assume a flat continuum in . We then treat the underlying continuum flux density and the total H flux as the only free parameters and determine them by fitting the multi-band photometry as described below.
We model the observed flux densities in the F356W, F410M, F444W, and F480M bands as the sum of the flat continuum and the integrated fluxes of the relevant emission lines falling within each filter at . For F356W, the model includes H and the [Oāiii]Ā 4960,5008 lines; for F444W and F480M, only H; and for F410M, no strong emission lines. The relative intensities of the [Oāiii] doublet are fixed to the theoretical ratio (), and H is tied to H through the assumed Balmer decrement. We then calculate the synthetic photometry for the modeled spectrum composed of the underlying continuum and H, H, and [Oāiii] lines and compare them with the NIRCam photometry. The latest NIRCam photometry for GLIMPSE-16043 is presented in the Appendix A. Based on the minimization approach, we obtain the best-fit continuum flux density of nJy and H line flux of . This best-fit H line flux estimate is in excellent agreement with the direct G395M measurement (Section 3.2), providing strong validation for our absolute flux calibration, the emission-line measurements, and the resulting continuum estimate.
In FigureĀ 2, we present the best-fit model at 3ā5m composed of the underlying continuum and emission lines (red curve), together with the corresponding synthetic photometry in the F356W, F410M, F444W, and F480M filters (red square). The model-predicted photometry agrees with the observed NIRCam fluxes mostly within the uncertainties, supporting the validity of our continuum and line estimates. This result demonstrates that the observed broadband fluxes can be reproduced by a combination of an exceptionally faint continuum and strong emission lines ā including [Oāiii] ā without requiring the purely PopĀ III-dominated SED proposed in fujimoto2025.
In FigureĀ 3, we explore the degeneracy between the continuum level and the assumed [Oāiii]/H ratio in reproducing the observed photometry of GLIMPSE-16043. In this test, we repeat the continuum-fitting procedure described above while fixing [Oāiii]/H to values between 0 and 8 in increments of 0.1. For each fixed ratio, we perform the same minimization over the continuum and H flux, and we color-code the results by the resulting value, computed using the synthetic photometry from the best-fit model and the observed F356W, F410M, F444W, and F480M fluxes.
Overall, the best-fit continuum level decreases as [Oāiii]/H increases, reflecting the larger fraction of the F356W excess that can be attributed to emission lines. However, for [Oāiii]/H, the reduced H and H fluxes require a higher continuum level to match the F444W and F480M photometry. The regime with [Oāiii]/H is disfavored by a rapid increase in , but the range [Oāiii]/H remains highly degenerate if the continuum level is not directly detected. This result cautions against applying photometric PopĀ III selection methods when the rest-frame optical continuum is undetected in all relevant filters.
As discussed in scenarioĀ (iv) of SectionĀ 4.1, however, the faint-continuum solution leads to another set of extreme physical properties: an unusually strong Balmer jump and very large EW(H). In fact, our best-fit values imply EW(H) Ā Ć and a Balmer jump strength, quantified by the difference in AB magnitudes between F200W () and the best-fit optical continuum () of Ā mag. Albeit the large uncertainty, the EW(H) exceeds the maximum values predicted by stellar+nebular photoionization models even under very top-heavy initial mass functions and/or the presence of very massive stars () (see e.g., TableĀ 2 in schaerer2024). Likewise, the inferred Balmer jump is nearly twice as large as the values reported in other high-redshift galaxies (e.g., cameron2024; katz2024).
This combination of extreme EW(H) and Balmer jump explains exactly why the fujimoto2025 analysis favored an [Oāiii]/H solution within the ā range, despite similar values from the simple parametric fits (FigureĀ 3), although the large parameter space of metal-enriched galaxies were explored. In the next section, we investigate possible models that can reproduce the remarkably strong EW(H) and Balmer jump observed in GLIMPSE-16043
4.3 Modeling GLIMPSE-16043
To investigate whether the observed strength of the H line and Balmer jump in GLIMPSE-16043 can be explained within standard photoionization frameworks, we perform a series of modeling experiments considering both stellar+nebular synthesis and pure nebular emission scenarios. Each model is redshifted to , and we compute NIRCam/F200W synthetic photometry and the continuum level at 4.1 m to compute the Balmer jump amplitude of from the model, for the sake of fair comparison with observations of GLIMPSE-16043.
(A) BPASS + Cloudy (stellar+nebular).
We first adopt binary stellar populations from BPASS (including the most metal-poor sets) as ionizing sources, coupled with Cloudy nebular models spanning and . These models reproduce the qualitative trends of increasing EW(H) and Balmer-jump amplitude with younger ages and higher ionization. However, the maximum EW(H) and Balmer-jump amplitude reach only āĆ and mag, respectively, falling short of the observed values by .
(B) PopĀ III + Cloudy.
PopĀ III ionizing spectra from zackrisson2011, coupled to the same nebular grids, behave similarly to (A) in their Balmer-jump predictions, with two notable differences. At very young ages ( Myr), these models can produce extremely large H equivalent widths up to āĆ , consistent with the measured EW range. However, even in these extreme cases, the Balmer jump remains far too weak. Moreover, PopĀ III-like systems are inconsistent with the observed [Oāiii]/H ratio, although PopĀ III stars may reside in metal-enriched gas during a āself-pollutedā phase following the first supernovae (fwang2012; rusta2025).
(C) Pure Nebular.
To test whether nebular continua alone can account for the large Balmer jump, we compute the nebular continuum emission, composed of freeābound, freeāfree, and two-photon emission, on grids of and using PyNeb. The models show stronger Balmer jumps with decreasing and increasing , although the latter also produces redder rest-frame UV and optical continua. No combination of parameters yields the observed Balmer jump: the maximum amplitude reaches only mag, even under extreme assumptions suppressing two-photon, freeāfree, and helium continuum contributions. Furthermore, the resulting red UV continua are inconsistent with GLIMPSE-16043, which exhibits a much bluer slope of .
(D) Blackbody + Cloudy.
The pure-nebular results suggest that an additional component in the rest-frame UV may be required to reproduce both the Balmer jump amplitude and the observed UV continuum shape. We therefore experiment with idealized blackbody (BB) sources as ionizing inputs, peaking in the rest-frame UV (ā5.2). Such blackbodies with ā5 are observed in tidal-disruption event (TDE) sources in the local Universe (gezari2021), making this scenario not entirely unfeasible. These models behave similarly to the BPASS and PopĀ III runs at very young ages ( Myr): hotter BBs ( K) enhance [Oāiii] but weaken the Balmer jump, while cooler BBs ( K) yield colors closer to the observations but still fail to reproduce the extreme Balmer-jump amplitude, reaching at most mag.
(E) BPASS/PopĀ III/BB + Cloudy with Fixed .
Finally, we also run Cloudy models for all three ionizing sources (BPASS, PopĀ III, BB) with the electron temperature fixed rather than solved self-consistently. While the models above adopt relatively simple assumptions, more complex conditions may exist in reality (e.g., a distribution of ISM properties, non-uniform geometries, or variations in the IMF). The āfixed experiments therefore provide a flexible test to probe a broader parameter space and to assess whether such effects could alleviate the discrepancies with the data. Among these runs, the BB+Cloudy model with K and K comes closest to the observed locus, yielding a Balmer-jump amplitude of mag. Although still insufficient, this value lies within of the measurement. However, the corresponding [Oāiii]/H remains significantly lower than the observed ratio of . While such a discrepancy might in principle be mitigated in a two-zone ISM configuration, where strong [Oāiii] emission arises from compact young star-forming regions, our results indicate that even the flexible setups of (E) cannot simultaneously reproduce all of the observed properties of GLIMPSE-16043 without tension.
| Model | Significant Balmer jump | Strong EW(H) | [Oāiii]/H | Blue rest-UV continuum |
| Ā mag | āĆ | |||
| (A) BPASS + Cloudy | () | () | ||
| (B) PopĀ III + Cloudy | () | () | ||
| (C) Pure nebular | () | () | ||
| (D) Blackbody (BB) + Cloudy | () | () | ||
| (E) BPASS/PopIII/BB + Cloudy (fixed ) | () | () | () |
TableĀ 2 summarizes these outcomes, highlighting the persistent tension between the models and the observed properties of GLIMPSE-16043. FigureĀ 4 also summarize the parameter spaces (left panel) and the closest SED (right panel) to the observed Balmer jump and the EW(H) properties. Overall, none of the models considered here are able to reproduce the observed properties of GLIMPSE-16043, in particular its unusually strong Balmer jump. This suggests that more exotic physical mechanisms may be at play in GLIMPSE-16043.
Note that we do not directly observe the Balmer jump in the G395M spectrum, and in principle alternative scenarios may remain, in which a very blue continuum extends smoothly from the rest-UV into the rest-optical, with the observed emission lines simply superimposed, eliminating the need for any significant discontinuity at the Balmer edge. To assess this possibility, we additionally fit the NIRCam photometry with a simple power-law continuum combined with the emission-line fluxes measured from the G395M data. We find that matching the extremely faint rest-optical continuum would require a very steep spectral slope (). Although extremely UV-blue SEDs can arise from young stellar populations with high escape fractions and a similarly blue galaxy has been reported at (; yanagisawa2025), GLIMPSE-16043 exhibits strong nebular emission lines, making a simple stellar origin difficult to reconcile with the this power-law SED scenario. This may motivate us to explore other SEDs, such as accretion disk models. However, the required slope of is significantly bluer than can be produced by standard geometrically thin, optically thick accretion disks, whose asymptotic slope is limited to (e.g., fabian2025). Moreover, the observed rest-UV photometry of GLIMPSE-16043 is inconsistent with any single power-law continuum. Consequently, such a smooth, jump-free continuum model is even less plausible than any of the physically motivated scenarios explored in cases (A)ā(E). A detailed discussion of these power-law tests is provided in AppendixĀ B.
4.4 Possible Scenarios
Although we find the difficulty with any model to reproduce the observed properties without any tensions in SectionĀ 4.3, the results still indicate that the āfixed BB+Cloudy model with Ā K provides the closest match to GLIMPSE-16043, especially the significant Balmer jump feature. Here we discuss possible physical scenarios that could produce such a hot, blackbody-like continuum component.
(i) Tidal-disruption event (TDE).
A first possibility is that we are witnessing a TDE, in which a star passing too close to a massive black hole is torn apart by tidal forces. At , the measured of GLIMPSE-16043 (fujimoto2025) corresponds to a rest-frame erg s-1 Hz-1, or an integrated UV luminosity of erg s-1. Adopting a bolometric correction of a factor of a few (appropriate for thermal TDE spectra; e.g., gezari2021), the implied bolometric luminosity is erg s-1. This is 1ā2 orders of magnitude fainter than the brightest local TDEs observed in the optical/UV ( erg s-1), but falls well within the observed distribution of lower-luminosity events (vanvelzen2020). It is also possible that the JWST observations captured the source during a fading phase, after the initial peak luminosity had already subsided, which would naturally make the event appear fainter. Interestingly, GLIMPSE-16043 shows a point-source morphology with an upper limit on the effective size of Ā pc (fujimoto2025). Such a high stellar surface density can promote the formation of an intermediate-mass black hole (IMBH) through runaway stellar collisions, and subsequently induce TDEs of stars (rantala2025). Therefore, the compact nature of GLIMPSE-16043 is in line with the TDE scenario.
We note that reproducing the observed Balmer jump requires the single blackbody component to peak tightly in the rest-frame UV. If the peak were shifted to longer wavelengths, the Balmer jump would be suppressed due to the contributions from the tail of the blackbody, while shifting to shorter wavelengths would make the red nebular continuum appear in the rest-frame UV regime. The best-fitting case with therefore represents a rather finely tuned solution. Remarkably, this temperature is in excellent agreement with the effective temperatures commonly observed in local TDEs, providing additional support for the TDE interpretation.
We further examined archival HST/WFC3 imaging taken in Oct.Ā 2015āApr.Ā 2016 (lotz2017). The GLIMPSE observations were obtained in Sep.Ā 2024, corresponding to a temporal separation of Ā yr in the observer frame (Ā yr in the rest frame). At the location of GLIMPSE-16043, we measure the flux density of nJy in HST/F160W, compared to nJy in NIRCam/F150W. Although the HST data are too shallow to draw firm conclusions, the non-detection in HST is consistent with the low luminosities implied by the JWST observations and does not contradict the TDE scenario.
(ii) Microquasar.
An alternative explanation is that the observed emission arises from a microquasar-like system, extending into the āminiquasarā regime (e.g., khlen2005), centered around an intermediate-mass black hole (IMBH). Microquasars are X-ray binaries in which a stellar-mass black hole accretes gas from a companion star, producing a luminous accretion disk and relativistic jets that resemble a scaled-down version of quasar activity. In such stellar-mass black hole X-ray binaries, the intrinsic thin disk peaks in soft X-rays (Ā keV) and is well described by a multi-temperature disk rather than a single blackbody (e.g., mcclintock2006). Several of the physical mechanisms operating in these systems, including a radiatively efficient inner disk, high mass-accretion rates that generate a luminous disk corona, and the presence of dense, scattering-dominated outflows driven by super-Eddington accretion, can also occur in the case of an IMBH accreting at super-critical rates. These conditions can substantially boost the luminosity and produce an emergent UV/optical continuum dominated by reprocessed emission, effectively extending the classical microquasar phenomenology into the āminiquasarā regime. Moreover, in super-Eddington accretion states with optically thick disk winds (e.g., SSĀ 433; Margon1984), the reprocessed emission from the outflow can produce a quasi-thermal optical/UV component that is approximately blackbody-like with āĀ K (fabrika2004; poutanen2007). This temperature range coincides with the best-fit BB+ model for GLIMPSE-16043, suggesting that a microquasar-like source embedded in a dense gaseous envelope could also plausibly account for the observed Balmer jump. In this interpretation, the extreme equivalent width of H would then arise from the combination of strong nebular recombination in the surrounding medium and a very weak underlying stellar continuum.
In summary, both the TDE and microquasar scenarios can, in principle, supply the hot, blackbody-like continuum required to reproduce the significant Balmer jump SED shape observed in GLIMPSE-16043. While the available data do not allow us to distinguish between them, the faint TDE and the microquasar disk winds both provide plausible pathways. Future deep time-domain monitoring will be critical to test these scenarios.
5 Revisiting Photometric PopĀ III Search
Our and recent spectroscopic results motivate a reassessment of the photometric selection of PopĀ III galaxy candidates. First, our NIRSpec follow-up of GLIMPSE-16043 has demonstrated that this source, originally identified by fujimoto2025 using the color- and SED-based selections, is not a bona-fide PopĀ III system but instead an exotic galaxy with an extraordinarily strong Balmer jump. This case demonstrates that a subset of the original color space is vulnerable to contamination by such extreme Balmer-jump objects. Second, a new spectroscopic PopĀ III candidate at has recently been reported in the Abell2744 field, AMORE6 (morishita2025). This source lies within the UNCOVER footprint that was photometrically searched by fujimoto2025. However, it turns out that AMORE6 is not included in the UNCOVER+Mega DR3 catalog used in that analysis due to blending with nearby bright galaxies, remarking the caution for the possible incompleteness due to the chance projection with bright foreground galaxies.
Using a -diameter aperture photometry for AMORE6-B on the NIRCam images with carefully subtractions of the foreground galaxies provided by morishita2025111The photometry is newly added in the revised version of morishita2025, which is shared by T.Ā Morishita in private communication. , we find that its photometry places the source well within the PopĀ III color selection in all the three diagnostic planes originally presented in fujimoto2025. We also find that its SED-based properties also satisfy the selection function, suggesting that AMORE6 would have been identified as a promising candidate. These results emphasize the importance of ensuring that future photometric search are both robust against exotic contaminants, like GLIMPSE-16043, and sufficiently inclusive so as not to miss promising PopĀ III candidates like AMORE6.
Guided by these two lessons, in the following subsections we refine the photometric color selection (Sec.Ā 5.1), reapply the revised criteria to the legacy survey fields originally analyzed by fujimoto2025 and five new lensing cluster fields publicly released by the CANUCS team (sarrouh2025) (Sec.Ā 5.2), and update the inferred PopĀ III UV luminosity function (UVLF) and PopĀ III cosmic star formation rate density (CSFRD) (Sec.Ā 5.3).
5.1 Refining Selection
FigureĀ 5 shows the original colorācolor diagrams presented in fujimoto2025, along with the positions of GLIMPSE-16043 and AMORE6. The fiducial PopĀ III selection was defined by the black dashed regions and designed to maximize purity by isolating sources with strong H and Balmer-jump signatures while minimizing [Oāiii] contamination. Our spectroscopic follow-up, however, has demonstrated that exotic Balmer-jump galaxies such as GLIMPSE-16043 can also fall within this space, producing a clear contamination pathway. To illustrate this, the black curve in FigureĀ 5 traces the locus occupied by such systems. It is constructed in the same manner as in FigureĀ 3: we vary the [Oāiii]/H ratio over the range 1.0ā2.5, where the best-fit continuum becomes extremely faint and therefore produces a strongly Balmer-jumpālike SED. For each value, we compute the corresponding continuum level together with the H (+H and [Oāiii]) emission, and then convolve the resulting spectra with the NIRCam filter response curves to generate synthetic photometry.
To address these issues, we revise the photometric selection by explicitly excluding the region traced by extreme Balmer-jump objects. The updated selection windows, shown by the red dashed lines in FigureĀ 6, remain consistent with the simulated PopĀ III color distributions (cyan points) while avoiding the locus of Balmer-jump contaminants. Formally, the revised selection is defined by the following polygons:
ā¢
F277W F356W vs. F410M F444W:
(1) -
ā¢
F200W F410M vs. F356W F444W:
(2) -
ā¢
F277W F356W vs. F410M F444W:
With these revised cuts, we confirm that AMORE6 (red filled pentagon) still falls well within the PopĀ III regions, demonstrating that inclusion is preserved in the updated definition.
In fujimoto2025, the ultimate candidate was obtained by combining the color-based selection with the SED-based criterion. The latter was evaluated using EAZY (brammer2008) with both metal-enriched and PopĀ III template sets, and comparing the quality of the fits. Specifically, defining the chi-square values from the two template sets as and , and , the selection required
or
However, the photometry may benefit from exceptionally high S/N owing to its strong lensing magnification. In such cases, even the best-fit PopĀ III models yield high absolute chi-square values, while the relative preference could remain overwhelming. To accommodate this situation, we extend the original criteria by adding an additional alternative requirement: in addition to Equations 4 and 5, a source is also accepted if
thus the SED-based criterion is satisfied if any of Equations 4, 5, or 6 holds. This ensures that highly magnified, high-S/N sources are not missed despite their intrinsically elevated values.
In addition, motivated by the lesson from GLIMPSE-16043 that emphasizes the importance of directly detecting the underlying rest-frame optical continuum, we impose an additional requirement,
ensuring that candidates are detected in the F410M band and validity of the flux excess measurements in F356W and F444W. Furthermore, because redshift reliability is a key component of the selection, we also require that the PopĀ III-based photometric redshift from EAZY () lies within ā6.6 with a uncertainty of , and that the source remains undetected in the optical with . Together, these additional criteria restrict the selection to sources with consistent redshifts at ā6.6.
We re-evaluated the contamination rate and completeness of these updated color- and SED-based selections following the same Monte Carlo approach as in fujimoto2025, generating mock photometry from simulated PopĀ III SEDs (for completeness) and metal-enriched galaxies (for contamination rate) with observational uncertainties applied. In FigureĀ 8, the resulting completeness and contamination rate distributions demonstrate that while the new selection is more restrictive, it retains high recovery efficiency for PopĀ III-like SEDs while substantially reducing the contamination risk from exotic Balmer-jump galaxies.
5.2 Re-analysis for Legacy and New Fields
We next re-apply the updated color- and SED-based selections to the legacy survey fields initially analyzed in fujimoto2025 as well as to newly available datasets. In fujimoto2025, the photometric search in the AbellĀ 2744 field was based on the UNCOVER+Mega DR3 catalog. However, as discussed in SectionĀ 5.1, this catalog missed the promising PopĀ III candidate AMORE6 due to blending with bright cluster galaxies. We therefore adopt the independent ALT catalog (naidu2024), which employs more aggressive source segmentation and successfully recovers AMORE6. For AMORE6, however, we use the dedicated aperture photometry measured on the NIRCam images with careful subtraction of the nearby foreground galaxies, provided by morishita2025, given the possible flux contamination from the nearby galaxies. For the other fields analyzed in fujimoto2025, we employed the same photometric catalogs as in the original study, namely: GLIMPSE (e.g., atek2025; see also kokorev2025), CEERS (e.g., finkelstein2024), PRIMER-UDS and PRIMER-COSMOS (e.g., donnan2024), GOODS-S, including JADES (e.g., eisenstein2023a; eisenstein2023b), JADES Origin Fields (JOF; robertson2023), and FRESCO (e.g., oesch2023). This ensures consistency with the earlier analysis.
In addition to revisiting the UNCOVER field, we also exploited the recently released CANUCS/Technicolor DR1 catalogs and lens models (sarrouh2025), which provide deep, homogeneous JWST/NIRCam imaging of five massive lensing clusters (AbellĀ 370, MACS0416, MACS0417, MACS1149, and MACS1423). These fields offer excellent depth (29ā30Ā mag) and lensing magnification, enabling sensitive searches for ultra-faint PopĀ III candidates in more survey areas. We apply the revised color cuts (FigureĀ 6) and the updated SED-based criteria (EquationsĀ 4ā7) uniformly across the ALT, CANUCS, and the other legacy catalogs.
From this re-analysis, we find that no additional sources satisfy all of our selection criteria as robust PopĀ III candidates beyond AMORE6. However, we identify four tentative candidates that meet most of the requirements but fail either one of the updated colorācolor criteria or the new S/N threshold in F410M (Eq.Ā 7). Interestingly, two of these four tentative candidates reside in known strongly lensed () LAEs at and in AbellĀ 370 (ade2022). The nature of these strong LAEs is reminiscent of recently reported spectroscopic PopĀ III candidates at ā6 (cai2025; morishita2025; vanzella2025). Moreover, one of the two is identified as a compact stellar clump within a strongly lensed arc, where other regions in the arc exhibit strong [Oāiii] emission, suggesting the presence of a metal-poor pocket embedded within an otherwise enriched galaxy and highlighting the power of spatially resolved studies enabled by strong lensing. Although we do not include these tentative sources in the following analyses, their colors are shown in FigureĀ 6 for reference, and further details for the spec- confirmed, two candidates identified in LAEs are provided in AppendixĀ C. We note that another tentative candidate initially reported in fujimoto2025, JOF-21739, fails two of the three updated colorācolor criteria and is therefore even less prominent PopĀ III candidate compared to the other tentative candidates.
For AMORE6, the colors are shown in FigureĀ 6, and the best-fit SED is presented in FigureĀ 7. Although deep spectroscopic follow-up is required to confirm the complete absence of the metal emission lines for for AMORE6, this outcome highlights the effectiveness of the refined selection in excluding contaminants such as exotic Balmer-jump galaxies while preserving promising PopĀ III candidates.
5.3 Updated PopĀ III UVLF and CSFRD
Given the recent example of LAP1-B, initially identified as a PopĀ III candidate due to the apparent absence of [Oāiii] in its NIRSpec spectrum (vanzella2023), but later revealed to exhibit multiple metal lines in deeper spectroscopy (nakajima2025), deep follow-up observations will be essential to first definitively establish whether AMORE6 is truly a PopĀ III system. Nevertheless, for the purpose of exploring the observational implications, we proceed under the assumption that AMORE6 is a PopĀ III galaxy and revisit the resulting constraints on the PopĀ III UV luminosity function (UVLF) and cosmic star-formation rate density (CSFRD) at ā7. We adopt the same methodology as in fujimoto2025, combining the magnification correction for the survey volume, completeness correction, and Poisson statistics to derive the volume density of the PopĀ III galaxies. Our reanalysis expands the survey volume by including the five CANUCS cluster fields (sarrouh2025) in addition to the legacy datasets. We summarize the effective survey areas in AppendixĀ D.
FigureĀ 9 presents the updated PopĀ III UVLF. The red filled pentagon denotes the spectroscopic candidate of AMORE6, and the red arrows mark the upper limits in the other magnitude bins. For the spec- candidates, the constraint remains broadly consistent with theoretical expectations from recent simulations (visbal2020; liu2020; sarmento2022; venditti2023; venditti2025). The corresponding number densities in each magnitude bin are summarized in TableĀ 3.
| ā [10-4 cMpc-3 dex-1] | ||
| 0 | ||
| 0 | ||
| 0 | ||
| 0 | ||
| 0 | ||
| 1 | ||
| 0 |
| Redshift | log() [Mā yr-1 Mpc-3] |
| 5.6ā6.6 | [ : ] |
We also derive constraints on the PopĀ III CSFRD by converting the UV luminosity density into a star-formation rate density using a PopĀ III-specific UV conversion factor (e.g., schaerer2002). We derive the lower and upper limits in the same manner as fujimoto2025: the lower limit is set by the detection of AMORE6 alone, while the upper limit is obtained by rescaling the general galaxy UVLF to match the upper boundary of the data point by AMORE6 and integrating it down to .
FigureĀ 10 presents the updated CSFRD constraints, which are slightly lower than those in fujimoto2025, reflecting the larger survey volume but the absence of additional detections. For comparison, we also show theoretical predictions (colored curves; maio2011; pallottini2014; xu2016; jaacks2019; sarmento2022; skinner2020; visbal2020; liu2020; munoz2022; hegde2023; venditti2023). Our updated limits remain broadly consistent with several theoretical predictions (e.g., jaacks2019; liu2020; hegde2023). They are also generally in line with the previous upper-limit estimate at ā (; nagao2008), which was based on searches for dual Ly and [Heāii] emitters, although the level of incompleteness in that approach remains uncertain. Our updated limits on the PopĀ III SFRD are summarized in TableĀ 4.
We caution that our current search strategy is optimized for galaxies and star clusters dominated by PopĀ III stellar populations, and may miss more complex systems. In particular, systems that have already been partially enriched by the first supernovae but still host surviving PopĀ III star clusters are likely difficult to identify with our criteria (e.g., fwang2012; rusta2025). Similarly, galaxies with a low nebular covering fraction or not very young PopĀ III star clusters (Ā Myrs) may fail to exhibit the very strong H features required by our selection, and would thus be underrepresented. The first case may be revealed by spatially resolved studies of strongly lensed arcs, such as LAP1 (vanzella2023; nakajima2025) and one of our tentative candidates found as a small clump in the strongly lensed arc (see Appendix C), which might represent metal-free pockets inside enriched systems. The ongoing large JWST program VENUS (PIDĀ 6882; PI: S.Ā Fujimoto & D.Ā Coe), which will establish homogeneous NIRCam 10-filter imaging for 60 massive lensing clusters, is expected to accelerate the discovery of such metal-free pockets and AMORE6-like pure PopĀ III systems. Regardless, future simulations that incorporate these observational constraints and adopt a self-consistent definition of PopĀ III galaxies will be essential to fully interpret the observational limits and to trace the cosmic history of PopĀ III star formation.
6 Summary
In this paper, we present deep JWST/NIRSpec G395M spectroscopy of GLIMPSE-16043, originally a promising PopĀ III candidate reporeted in fujimoto2025 through JWST/NIRCam photometry. Our main findings are as follows:
1.
We performed 30.4Ā hr of on-source NIRSpec/G395M observations. The spectra clearly detect H, H, and [Oāiii], with line ratios of H/H and [Oāiii]/H. These measurements show that GLIMPSE-16043 is not a genuine zero-metallicity PopĀ III system.
2.
By combining the line fluxes with NIRCam photometry, we constrain the underlying continuum and find an extraordinarily strong Balmer jump ( mag) and an extreme rest-frame H EW (āĆ ). These values exceed predictions from all current stellar+nebular and pure nebular photoionization models.
3.
We test a suite of models: (A) BPASS+Cloudy and (B) PopĀ III+Cloudy reproduce some qualitative trends with the strong EW(H) but fail to match the Balmer jump; (C) pure nebular continua cannot simultaneously fit the Balmer jump and UV slope; (D) blackbody+Cloudy models with K enhance [Oāiii] but weaken the Balmer jump; and (E) blackbody+Cloudy models with fixed provide the closest match with a K blackbody and K, yielding a Balmer jump amplitude of mag, still insufficient to reproduce all observed properties. This points toward more exotic origins.
4.
The origin of the singleāblackbody, peaking in the rest-frame UV, may plausibly arise from either (i) a tidal-disruption event with Ā ergĀ s-1, or (ii) a microquasar-like system in which super-Eddington disk winds generate a pseudo-photosphere at āĀ K. Both scenarios can, in principle, account for the best-fit blackbody component.
5.
These results demonstrate that photometric PopĀ III color criteria are vulnerable to contamination by exotic Balmer-jump galaxies, especially when the rest-frame optical continuum is undetected. To mitigate this, we refine the fujimoto2025 selection by excluding the locus traced by such objects. The revised criteria also recover the recently reported spectroscopic candidate AMORE6 (morishita2025), while excluding GLIMPSE-16043. We emphasize, however, that the absence of AMORE6 in the earlier search of fujimoto2025 was due to blending with nearby bright foreground galaxies, not a failure of the selection itself. This indicates that both the original and updated criteria maintain reasonable completeness, while the refined selection more effectively removes key contaminants.
6.
Reapplying the refined criteria to the JWST legacy fields (UNCOVER, GLIMPSE, CEERS, PRIMER, JOF, JADES, FRESCO) and to the newly publicly released CANUCS cluster fields, we do not identify any additional sources that satisfy our PopĀ III selection, beyond AMORE6. Still, we find four tentative candidates that meet most of the requirements but fail either one of the updated colorācolor criteria or the new S/N threshold in F410M. Interestingly, two of these tentative candidates found in known strongly lensed LAEs at , reminiscent of recently reported spectroscopic PopĀ III candidates at ā6 (cai2025; morishita2025; vanzella2025).
7.
Based on this re-analysis, and adopting AMORE6 as a PopĀ III galaxy for the purpose of exploring the observational implications, we update the PopĀ III UV luminosity function and cosmic star-formation rate density at ā7. The larger survey area leads to constraints slightly lower than those reported in fujimoto2025, yielding ā, which falls within the range of current theoretical predictions. However, these values should be regarded as lower limits, since our selection is optimized for pure, very young, highānebular-covering-fraction PopĀ III systems and is likely incomplete for more complex or partially enriched galaxies that may still host surviving PopĀ III star clusters. A self-consistent definition of PopĀ III populations, coupled with simulations that incorporate such observational selection effects, will be essential for fully interpreting the constraints on PopĀ III star formation across cosmic time.
Appendix A Updated photometry of GLIMPSE-16043
The photometry used in fujimoto2025 was based on an early version of the data reduction and an internal catalog, whereas in this study we adopt the latest photometry for GLIMPSE-16043. All updated fluxes remain within the uncertainties of the original measurements and do not alter the conclusions of fujimoto2025. In fact, the new photometry even strengthens the previously reported photometric PopĀ III signature: the quantity () increases when using the updated measurements. A comprehensive description of the revised imaging reduction and catalog construction is provided in the GLIMPSE survey release paper (atek2025); here we summarize the key updates relevant to this work.
Although the procedures for source detection, artifact removal, bCG subtraction, and catalog flagging follow the same principles as those used during the initial identification of GLIMPSE-16043 (fujimoto2025), a major update is the adoption of custom flat-fields. Specifically, we generated improved flat-field calibration frames using all publicly available NIRCam observations as of January 12, 2025. This is done due to the fact that a set of standard STScI supplied flat-fields introduced substantial correlated noise into the final mosaics, detrimentally affecting the effective depth of the images and introducing numerous spurious sources, especially in regions with where few dithers overlap. This problem was further exacerbated by the GLIMPSE observing strategy itself, which employed six wide dithers supplemented by four sub-pixel dithers at each position to optimize PSF sampling. While this pattern is ideal for PSF reconstruction, it also compounds flat-field residuals during co-addition. These residuals can resemble high signal-to-noise sources, particularly toward the mosaic edges where fewer large dithers overlap (e.g. see the discussion in kokorev2025). Implementing our custom flat-fields mitigates these effects and yields a measurable improvement in depth, and therefore S/N, reaching up to 0.5 magnitudes in the long-wavelength filters and approximately 0.2 magnitudes in the short-wavelength channels.
| āāāR.A. [deg] | āāā342.2123718 |
| āāāDec. [deg] | āāā44.528751 |
| āāāF090W [nJy] | āāā |
| āāāF115W [nJy] | āāā |
| āāāF150W [nJy] | āāā |
| āāāF200W [nJy] | āāā |
| āāāF277W [nJy] | āāā |
| āāāF356W [nJy] | āāā |
| āāāF410M [nJy] | āāā |
| āāāF444W [nJy] | āāā |
| āāāF480M [nJy] | āāā |
Appendix B Power-law and emission-line model
In SectionĀ 4.3, we discuss that physically motivated stellar+nebular or pure photoionization models struggle to reproduce the observed SED properties of GLIMPSE-16043. However, because the G395M spectrum does not directly detect a Balmer jump, we also test an extreme ājump-freeā scenario in which a single, very blue power-law continuum extends smoothly from the rest-UV into the rest-optical wavelength. To assess this possibility, we fit the NIRCam photometry with a single power-law continuum,
combined with the H, H, and [Oāiii]ā4960,5008 fluxes measured from the G395M data. We set and the continuum normalization as free parameters.
In FigureĀ 11, we present the best-fit power-law + emission line model. The best-fit model requires an extremely steep slope of . Such extremely UV-blue SEDs can arise from young stellar populations with high escape fractions (e.g., yanagisawa2025), GLIMPSE-16043 exhibits strong nebular emission lines, making a simple stellar origin difficult to reconcile with the this power-law SED scenario. This motivates to other SED models, such as black hole accretion disk models (e.g., tacchella2025; ji2025; fabian2025), where the strong nebular emission may arise from surrounding narrow-line regions in the black hole accretion disk. Nevertheless, this is far bluer than physically achievable in standard, geometrically thin, optically thick accretion disks, whose power-law slope is analytically limited to (; e.g. shakura1973) and challenging to produce significantly bluer continua. Even with , the model still underpredicts the rest-frame optical continuum inferred from our G395M-based continuum estimate by at 4Ā m, implying that an even more extreme slope () would be required well outside the realm of any plausible accretion-disk or radiative reprocessing model. Furthermore, the observed rest-UV photometry of GLIMPSE-16043 is incompatible with such a simple power-law shape, and the implied rest-frame EW(H) becomes even more extreme, āĆ . These tests indicates that a smooth, jump-free continuum is even more challenging to reproduce the observed SED of GLIMPSE-16043 than the models discussed in SectionĀ 4.3.
Appendix C Tentative Candidates
In SectionĀ 5.2, our re-analysis in all public legacy JWST fields both in general and lensing cluster fields show that no additional sources satisfy all of our updated selection criteria as robust PopĀ III candidates at ā6.6, beyond AMORE6. However, we identify four tentative candidates that meet most of the requirements but fail either one of the updated colorācolor criteria or the new F410M S/N threshold (Eq.Ā 7). Interestingly, two of these four sources reside in known strongly lensed LAEs, reminiscent of recently reported spectroscopic PopĀ III candidates at ā6 (cai2025; morishita2025; vanzella2025). Below, we describe the two tentative candidates associated with the lensed LAEs. Their best-fit PopĀ III and metal-enriched galaxy SED models are shown in FigureĀ 12.
A370-z6LAE-1c ā is identified as a tentative PopIII candidate residing within a strongly lensed LAE, dubbed A370-z6LAE-1, at behind the AbellĀ 370 cluster, where it appears as one of five resolved stellar clumps with a lensing magnification of . This magnification implies an intrinsic absolute UV magnitude of , based on the observed F150W magnitude of 28.6Ā mag with a -diameter aperture. The source coordinates are (RA, Dec) = (39.9805416, ), with the source ID in the CANUCS DR1 catalog is 2123895 (sarrouh2025). A370-z6LAE-1c satisfies our updated PopĀ III selection, including the SED-based criteria shown in Figure12, except for the third colorācolor diagram ([F356WF410M] vs. [F410MF444W]) in FigureĀ 6. Its location in this diagram remains consistent with PopĀ III model tracks, but corresponds to intermediate stellar ages (ā30 Myr), where the strength of H emission begins to diminish. As a result, the color space becomes degenerate with that of metal-enriched galaxies. Therefore, although the position of A370-z6LAE-1c does not rule out a PopĀ III nature, it highlights the strong degeneracy between age and metallicity in cases where H is weak, which is precisely the reason why our third color cut excludes the region where PopĀ III SED tracks overlap with enriched-galaxy models. For this reason, we classify A370-z6LAE-1c as a tentative candidate.
The left panel of FigureĀ 13 shows A370-z6LAE-1 in the image plane (i.e., observed frame without lensing correction), where the five clumps (ā) are clearly resolved in the CANUCS NIRCam imaging (sarrouh2025), and the Ly emission detected with VLT/MUSE (ade2022) is spatially coincident with all of them. The right panel presents the source-plane reconstruction of A370-z6LAE-1, obtained using the latest publicly released CANUCS lens model for Abell370 (sarrouh2025). We adopt F356W for the reconstruction because this band provides the highest S/N for clump among the NIRCam long-wavelength (LW) filters. The host system exhibits an intrinsically elongated morphology spanning kpc, consisting of a bright core (clumps and ) and a tail hosting clumps , , and . Clump lies between the brightest core clumpĀ and the tail clumpĀ , with projected separations of 1Ā kpc from both. This configuration is consistent with clumpĀ being either a metal-poor pocket within the host galaxy of A370-z6LAE-1 or an infalling metal-poor satellite. Importantly, the kpc separation ( at ) from enriched neighbors demonstrates that PopIII-like signatures in compact metal-poor pockets can be easily diluted at the nominal NIRCam resolution in general-field surveys. This highlights the power of spatially resolved searches aided by strong lensing.
To further characterize the system, we fix the Ly redshift and flux and perform SED modeling with Bagpipes (carnall2018) for each of the five clumps; their best-fit SEDs are shown in FigureĀ 14. The Bagpipes fits indicate that A370-z6LAE-1c represents a low-mass () stellar cluster hosted by a galaxy, and they confirm that clumpĀ exhibits very weak [Oāiii] emission, in contrast to the strong [Oāiii] seen in most of the other clumps. The weak [Oāiii] is also independently supported by the latest deep F360M imaging in this field (#5890; PIs: C.Ā Withers & A.Ā Muzzin), which shows no enhancement relative to F356W. Conversely, the NIRCam F360M grism data obtained in the same region (PIDĀ 2883; PI: F.Ā Sun) show marginal positive flux at the expected [Oāiii] wavelength inferred from the Ly redshift (integrated S/N 3), suggesting the possibility of non-zero [Oāiii] emission in A370-z6LAE-1c. A key caution for both the imaging and grism measurements of A370-z6LAE-1c is the presence of a nearby bright cluster-member galaxy, which makes the results highly sensitive to the details of local-background subtraction. In fact, an independent reduction and analysis of the F360M grism data yield the possible [Oāiii] feature much weaker with S/N , highlighting the level of practical uncertainty introduced by bright neighboring sources. Taken together, these results reinforce the tentative nature of A370-z6LAE-1c as a PopĀ III candidate, while simultaneously demonstrating how strong gravitational lensing enables us to probe distinct stellar populations inside a galaxy.
A370-z6LAE-2 ā is a compact LAE at found behind the Abell370 cluster, with a lensing magnification of . Its observed F150W magnitude of 30.3Ā mag (1.7; measured in a -diameter aperture) implies an intrinsic absolute UV magnitude of or even fainter. The source coordinates are (RA, Dec) = (39.9814916, ), with the source ID in the CANUCS DR1 catalog is 2114669 (sarrouh2025). A NIRCam color composite of the system is shown in FigureĀ 15.
A370-z6LAE-2 satisfies all of our updated PopIII selection criteria in both the colorācolor diagrams and the SED-based tests, except for the requirement of (Eq.7). Its measured S/N in F410M is only 1.4, which naturally reflects the faintness of the object. As discussed in Sec.Ā 5.1, this specific F410M requirement is designed to exclude contaminants exhibiting a significant Balmer jump, such as the case encountered for GLIMPSE-16043. Although the current position of A370-z6LAE-2 in FigureĀ 6 lies well away from the region occupied by the exotic Balmer-jump contaminants, the low S/N in F410M permits its inferred location to scatter into that regime. On the other hand, A370-z6LAE-2 is relatively isolated, with no bright neighboring sources, implying that deeper F410M imaging alone, even without the need for spectroscopy, could robustly establish its color and elevate it to a promising PopĀ III candidate.
Appendix D Survey Area
To compute the effective survey volume for identifying PopĀ III candidates at ā6.6, we follow the methodology of fujimoto2023b; fujimoto2025, in which the source-plane area as a function of lensing magnification is evaluated for each cluster field. For the seven lensing clusters analyzed in this work ā five CANUCS clusters (AbellĀ 370, MACSĀ 0416, MACSĀ 0417, MACSĀ 1149, and MACSĀ 1423), AbellĀ 2744 from UNCOVER, and AbellĀ S1063 from GLIMPSE ā we use the latest publicly released lens models provided by the CANUCS (sarrouh2025), UNCOVER (furtak2023a), and GLIMPSE (atek2025) teams. Because our ā6.6 PopĀ III colorācolor selection requires detections in all of F200W, F277W, F356W, F410M, and F444W, we restrict the survey area to regions where the full set of NIRCam bands overlaps. To this end, we manually define areas outlining the footprint in which all five filters are available at the required depth.
For each cluster, we compute the source-plane area as a function of magnification , binned logarithmically from to . At each pixel inside the defined region, the magnification is obtained from the lens model (assuming a source redshift of ), and the corresponding source-plane contribution is computed as the image-plane pixel area divided by . Pixels are then accumulated into cumulative magnification bins, such that each bin records the total source-plane area with magnification exceeding a given threshold, . We then convert the ā relation into a relation between and the effective limiting magnitudes for each field. The limiting depths are obtained by scaling according to the magnification and by adopting the F200W sensitivities reported in the literature for the lensing fields analyzed in this study. For non-lensed general blank fields, we use the detection limits and survey areas compiled in fujimoto2025. The resulting ādepth relations for all fields are presented in FigureĀ 16. These relations are used directly to compute the magnification-corrected survey volume employed in our updated PopĀ III UVLF and CSFRD analysis (SectionĀ 5.3).